OpenStudy (anonymous):

if tan^-1x + tan^-1y + tan^-1z= pi or pi/2 then prove x+y+z=xyz or xy +yz+zx=1

OpenStudy (anonymous):

dude i got this one

OpenStudy (anonymous):



OpenStudy (anonymous):

tan inverse let me write it as arctan ( for ease)- so arctan x +arctan y =arctan(x+y/1-xy) arctan(x+y/1-xy) + arctan z = arctan (x+y+z - xyz/ 1- xy - zx - zy); if tan pi = 0; therefore numerator = 0 solve; you will get it; tan pie/2 = infinity, denominator = 0; sove it and you will get it;

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

what is arc tan

OpenStudy (anonymous):

tan inverse is also called as arctan da.

OpenStudy (anonymous):

what about other questions

OpenStudy (anonymous):

still trying and asking